OpenStudy (anonymous):

What is the exact value of cos 225° cos 75°?

OpenStudy (anonymous):

by using identity cos a cos b = (1/2) [cos (a - b) + cos (a + b)] cos225cos75=1/2[cos(225-75)+cos(225+75)] =1/2[cos150+cos300] =1/2[-0.86+0.5] =1/2[-0.36] =-0.18

OpenStudy (anonymous):

so (1-sqrt3)/4?

OpenStudy (anonymous):

thank you